1. A nuclear chemist receives an accidental radiation dose of 5.0 Gy from slow neutrons ( RBE= 4.0 ). What does she receive in rad, rem and J/kg.

2. In a diagnostic x-ray procedure, 5.00 × 10 raise to 10 photons are absorbed by a tissue with a mass of .600 kg. The xray wavelength is .200 nm.

A. What is the total energy absorbed by the tissue
B. What is the equivalent dose in rem?
1
Expert's answer
2019-04-14T07:53:13-0400

1

rad=0.01 Gy

1 gray (Gy) =100 rad

5.0 Gy=500 rad

1 rad=1 rem

5.0 Gy=500 rem

1 gray = 1 joule/kilogram

5.0 Gy=500 joule/kilogram

Answer: 500 rad; 500 rem; 500 joule/kilogram


2

A.

"E=N\frac {hc}{ \lambda} (1)"

where N=5.00 × 1010, h=6.6 J×s, c=3× 108 m/s, λ=.200 nm

Using (1) we got: E=4.97×10-5 J


B

The equivalent dose is given by formula


"H= Dw (2)"

where D is the absorbed dose, w is the radiation weighting factor (for x-ray w=1)

Absorbed dose is defined as


"D=\frac {ΔE}{Δm} (3)"

Using (3) we got: D=8.28×10-5 J/kg

The equivalent dose is equal to 8.28×10-7 rem


Answer:

4.97×10-5 J

8.28×10-7 rem